# Phase transitions in two-component Bose-Einstein condensates with Rabi frequency (I): The De Giorgi conjecture for the local problem in $\mathbb{R}^{3}$

### v2

@@ Abstract.
In this series of papers, we investigate coupled systems arising in the study of two-component Bose-Einstein condensates, and we establish classification results for solutions of De Giorgi conjecture type.
In the first paper of the series, we focus on the local problem of the form
$\begin{cases}\Delta u=u(u^{2}+v^{2}-1)+v(\alpha uv-\omega),\\ \Delta v=v(u^{2}+v^{2}-1)+u(\alpha uv-\omega),\end{cases}$
and prove that positive global solutions in $\mathbb{R}^{3}$ satisfying $\tfrac{\partial u}{\partial x_{3}}>0>\tfrac{\partial v}{\partial x_{3}}$ must be one-dimensional.
@@ 1. Introduction
In the 1970s, De Giorgi ([17]) proposed a celebrated conjecture concerning the symmetry of solutions to semilinear elliptic equations. More precisely, let $u$ be a bounded entire solution of the Allen-Cahn equation
@@
This conjecture is strongly related to the Bernstein problem for minimal surfaces, as the level sets of $u$ behave asymptotically like minimal hypersurfaces. The De Giorgi conjecture and its variants play a fundamental role in the study of phase transition models, geometric analysis, and the qualitative theory of partial differential equations. In particular, they provide a bridge between elliptic PDEs, minimal surface theory, and problems arising in mathematical physics.
@@ 1.1. The system of Bose-Einstein condensates
In this paper, we extend the De Giorgi conjecture to a competitive system. We let $(u,v)$ be a positive pair satisfying the following system:
@@ 1.2. Classical theory in a special case $\alpha=2$
For such a single equation, there have been a number of important results related to the classification of global solutions. In [28], Ghoussoub and Gui proved the De Giorgi conjecture in $\mathbb{R}^{2}$, and then Ambrosio and Cabré [5] showed the same result in $\mathbb{R}^{3}$. In dimensions $\mathbb{R}^{4}-\mathbb{R}^{8}$, the validity of the De Giorgi conjecture remains open, except under additional assumptions. For example, Ghoussoub and Gui [29] proved the flatness of solutions under an additional odd-symmetry assumption. Besides, in [41], Savin proved the De Giorgi conjecture in $\mathbb{R}^{8}$, with the additional assumption that the zero set is an entire graph. In fact, the asymptotic flatness of the level set implies the flatness of a global solution.
These classification results have also been extended to more general settings. In [47, 50], the authors established the one-dimensional symmetry for solutions to the $p$-Laplacian Allen-Cahn equation. In [12, 22, 43, 42], the De Giorgi conjecture was verified for the fractional Allen-Cahn equation in various settings.
@@ 1.3. Gibbons’ conjecture: similarities and differences
Unlike the De Giorgi conjecture, for which the critical dimension is $8$, the validity of the Gibbons conjecture has been established in all dimensions by independent methods due to Barlow, Bass and Gui [7], Berestycki and Hamel [8], Berestycki, Hamel and Monneau [9], and Farina [23]. Farina’s contribution [23] was the first in this direction, though it required additional assumptions as noted in [9]. For the fractional case, we refer interested readers to [52, 53].
@@ 2. Main results and key ideas
then the pair $(u,v)$ is one-dimensional. More precisely, there exists a unit vector $\vec{\nu}\in\mathbb{R}^{3}$ and a pair of positive functions $(U(t),V(t))$ satisfying (1.2) in $\mathbb{R}$ such that

### v3

@@ Abstract.
We consider the following elliptic system modeling Rabi-coupled two-component Bose-Einstein condensates,
$\begin{cases}\Delta u=u(u^{2}+v^{2}-1)+v(\alpha uv-\omega),\\[2.84526pt] \Delta v=v(u^{2}+v^{2}-1)+u(\alpha uv-\omega),\end{cases}\qquad u,v>0\ \text{in }\mathbb{R}^{3},$
and establish the first De Giorgi-type classification result for this competitive system. We prove that any entire solution satisfying the monotonicity condition
$\partial_{x_{3}}u>0>\partial_{x_{3}}v$
is necessarily one-dimensional and connects the two stable equilibria determined by the parameters $(\alpha,\omega)$.
The proof relies on two new ingredients. First, a sharp range estimate reveals a strictly negative defect structure in the linearized operator, leading to a divergence-form identity that forces the directional slopes of all level sets to be constant. Second, we develop a translation-invariant energy evolution method for the associated coupled Ginzburg-Landau functional, yielding an $O(R^{2})$ energy-growth bound and a Liouville-type theorem.
Our results extend the classical classification results from the scalar Allen-Cahn equation to a physically relevant coupled system, and provide the first rigorous confirmation that De Giorgi’s phenomenon persists under Rabi coupling.
@@ 1. Introduction
In the 1970s, De Giorgi ([17]) proposed a celebrated conjecture concerning the symmetry of monotone solutions to semilinear elliptic equations. More precisely, let $u$ be a bounded entire solution of the Allen-Cahn equation
@@
This conjecture lies at the intersection of elliptic PDE, phase transition theory, and geometric analysis, since the level sets behave asymptotically like minimal surfaces, connecting the conjecture to the classical Bernstein problem. Over the years, it has inspired a vast literature and has become fundamental to the qualitative analysis of nonlinear elliptic equations, providing a natural bridge between elliptic PDEs, minimal surface theory, and problems in mathematical physics.
@@ 1.1. The system of Bose-Einstein condensates
The present paper extends the De Giorgi framework from single equations to a competitive elliptic system arising in two-component Bose-Einstein condensates (BECs). We consider positive solutions $(u,v)$ of
@@ 1.2. Classical theory in a special case $\alpha=2$
For such a single equation, there have been a number of important results related to the classification of global solutions. In [29], Ghoussoub and Gui proved the De Giorgi conjecture in $\mathbb{R}^{2}$, and then Ambrosio and Cabré [5] showed the same result in $\mathbb{R}^{3}$. In dimensions $\mathbb{R}^{4}-\mathbb{R}^{8}$, the validity of the De Giorgi conjecture remains open, except under additional assumptions. For example, Ghoussoub and Gui [30] proved the flatness of solutions under an additional odd-symmetry assumption. Besides, in [42], Savin proved the De Giorgi conjecture in $\mathbb{R}^{8}$, with the additional assumption that $u(x^{\prime},x_{n})\to\pm 1$ non-uniformly as $x_{n}\to\pm\infty$. In fact, the asymptotic flatness of the level set implies the flatness of a global solution. Later, the convergence assumption in [42] was replaced by the assumption that some level set of $u$ is an entire graph by Farina and Valdinoci in [26].
These classification results have also been extended to more general settings. In [48, 51], as well as the previously mentioned [26], the authors established the one-dimensional symmetry for solutions to the $p$-Laplacian Allen-Cahn equation. In [12, 22, 44, 43], the De Giorgi conjecture was verified for the fractional Allen-Cahn equation in various settings.
@@ 1.3. Gibbons’ conjecture: similarities and differences
Unlike the De Giorgi conjecture, for which the critical dimension is $8$, the validity of the Gibbons conjecture has been established in all dimensions by independent methods due to Barlow, Bass and Gui [7], Berestycki and Hamel [8], Berestycki, Hamel and Monneau [9], and Farina [23]. Farina’s contribution [23] was the first in this direction, and his work did not require the boundedness of the solution. For the fractional case, we refer interested readers to [53, 54].
@@ 2. Main results and key ideas
Remark 2.1.
In Farina-Sciunzi-Soave [24], the authors have already realized that the system (1.2) can be decoupled in the case $\omega=0$, and they have obtained a similar result as in Corollary 2.1.
@@
then the pair $(u,v)$ is one-dimensional. Precisely, there exists a unit vector $\vec{\nu}\in\mathbb{R}^{3}$ and a pair of positive functions $(U(t),V(t))$ satisfying (1.2) in $\mathbb{R}$ such that
@@
One of the key innovations of this paper is the combination of energy analysis with De Giorgi-type classification results, providing a deep understanding of the structure and stability of solutions in the coupled two-component Bose-Einstein condensate system. In particular, we introduce a novel method of proving the one-dimensional symmetry of solutions by constructing appropriate directional derivative symmetries and utilizing range estimates. This approach is crucial in analyzing the asymptotic behavior of the solutions and contributes to a deeper understanding of the system’s solution structure, especially in higher-dimensional settings.
Another key contribution lies in the introduction of an energy functional analysis, which serves as a powerful tool for handling the system’s stability and structural behavior. This approach not only provides a new perspective on the classical Ginzburg-Landau functional but also reveals the existence of nontrivial solutions in higher dimensions. Additionally, the classification of boundary conditions and steady-state solutions, which was previously underexplored in the literature, has been thoroughly analyzed in this study. Through detailed analysis of steady-state solution structures and convergence behavior, we unveil multiple solution structures under different parameter regimes, extending the classical Liouville-type theory to coupled systems.
From a methodological perspective, although De Giorgi-type classification results have been well-developed for single equations, handling coupled systems presents significant challenges due to the nonlinearity and interaction between components. Our breakthrough lies in combining symmetry-breaking and asymptotic analysis methods, which allow us to analyze the behavior of the solutions and prove their one-dimensional symmetry in $\mathbb{R}^{3}$. This method is particularly valuable for systems with complex interactions, such as Bose-Einstein condensates, and extends the applicability of De Giorgi-type classification results to higher-dimensional multi-component systems.
This work not only provides a classification of solutions for coupled systems but also opens new avenues for analyzing the behavior of solutions in high-dimensional spaces, which has not been fully explored in the literature. Furthermore, our findings offer a solid theoretical foundation for future studies of multi-component systems with nonlocal effects or interactions, particularly in the context of mathematical physics and the theory of phase transitions.